OpenStudy (anonymous):

tom savings and loan is paying 6% interest compounded monthly. how much will $100 deposited at the end of each month be worth after 2 years?

OpenStudy (tkhunny):

i = 0.06 -- Annual Interest Rate j = i/12 -- Monthly Interest Rate r = 1+j -- Monthly Accumulation Factor. Starting from the last payment \($100(1 + r + r^{2} + ... + r^{23})\) Because we know things about finite geometric series \($100\dfrac{1-r^{24}}{1-r} = $100\dfrac{r^{24} - 1}{j}\) Whether you remember the right formula or not, the method of Basic Principles can ALWAYS be followed and the correct formula constructed.